### 23 Apr 2026 v2

# Unitary Quadratic Quantum Gravity in 4D

Abstract
In quadratic gravity, with a positive Weyl squared coefficient, the extra spin–2 sector is shown to correspond to a dual inverted harmonic oscillator, instead of a ghost. Using the Wightman spectrum condition, we prove that the associated Källén–Lehmann spectral density vanishes, reflecting the absence of a normalizable ground state and the spacelike nature of the propagator pole. This uniquely fixes the propagator to a principal value form as a theorem, not a prescription. The optical theorem is satisfied, the dual IHO spin-2 is not an asymptotic state, and gives only virtual contributions at all loop orders. As a result, unitarity is preserved consistently with renormalizability.
I Introduction
Quadratic (Stelle) gravity [21],
$S=\int\!d^{4}x\sqrt{-g}\!\left[\frac{M_{p}^{2}}{2}R+\frac{\alpha}{2}R^{2}+\frac{\beta}{2}W_{\mu\nu\rho\sigma}W^{\mu\nu\rho\sigma}\right],$ (1)
is the unique perturbatively renormalizable local theory of quantum gravity in four spacetime dimensions [21, 20, 5]. This uniqueness is not an accidental technicality. In quantum field theory, perturbative renormalizability is the criterion that isolates predictive ultraviolet (UV) completions from merely effective low-energy descriptions, which is precisely why it underlies the success of the Standard Model [25]. For gravity, the higher derivative terms in (1) raise the UV falloff of the graviton propagator from $1/k^{2}$ to $1/k^{4}$, thereby yielding power counting renormalizability in four dimensions [21, 20]. The standard obstruction, however, is that the additional spin-2 degree of freedom has long been interpreted as a ghost, apparently spoiling unitarity [21, 20]. A variety of proposals have been developed to evade this difficulty, including fakeons [1], Lee–Wick-type prescriptions and related contour deformations [17, 6], and PT-symmetric quantization [3], all of which modify the interpretation of the extra pole while introducing structure beyond the minimal local theory.
In our companion work [16], we argued that this standard conclusion is not forced upon quadratic gravity. For the sign choice $\beta>0$, the extra spin-2 sector is not governed by a negative definite ghost Hamiltonian, but by a dual inverted harmonic oscillator (dual-IHO), namely an indefinite Hamiltonian system with hyperbolic phase space which shares close similarities with IHO. This distinction is decisive. A ghost is pathological because its negative definite Hamiltonian leads to negative norm excitations and a direct conflict with the probabilistic interpretation of quantum theory. By contrast, the IHO and dual-IHO are not negative definite systems (but indefinite); their essential quantum property is instead the absence of a normalizable ground state. The physical relevance of this structure is not peculiar to quadratic gravity. IHO dynamics is a universal template for controlled instability: it appears in the Berry–Keating $H=xp$ system [4], and in our broader framework it also underlies the quantization of the quadratic gravity spin-2 sector [16]. The point of the present paper is therefore not to introduce an ad hoc prescription, but to pinpoint the precise mechanism by which this universal IHO structure restores unitarity in quadratic gravity. The key issue is the propagator of the extra spin-2 field. For an ordinary particle degree of freedom, the full propagator is constrained by the Källén–Lehmann (KL) representation [11, 18], whose spectral density is defined on the timelike spectrum of physical states and is rooted in the Wightman spectrum condition of local quantum field theory [22, 8]. In the dual-IHO case, however, two facts hold simultaneously: the field admits no normalizable vacuum state, and its pole lies at spacelike momentum. These two facts are already sufficient to force the KL spectral density to vanish. Consequently, the propagator of the dual-IHO spin-2 field is fixed uniquely to its principal value form (PV). In other words, the principal value is not an externally imposed prescription; it is a theorem dictated by the Hilbert space structure and the spectral support of the theory [16].
In this paper, we sharpen that conclusion into a self-contained unitarity analysis with three main results. First, we show that the KL spectral density vanishes as a consequence of the Wightman spectrum condition applied to a field with spacelike momentum support, so that $\rho=0$ is not an assumption but a theorem. Second, we give a Cutkosky analysis about how the optical theorem is satisfied in our case, with an additional spin-2 field not corresponding to any on-shell states. Third, we perform the Landau analysis of the relevant bubble diagrams and compare it explicitly with the Feynman–Wheeler case of Ref. [2], thereby showing why the nonlocal divergences associated with timelike principal value propagators are absent when the pole is spacelike. The result is a minimal and sharp statement: in quadratic gravity with $\beta>0$, the extra spin-2 mode remains purely virtual, contributes to UV renormalization through its dispersive part, but never enters physical absorptive cuts. Unitarity and renormalizability, therefore, coexist within the original local theory [16].
We use the metric signature $\eta_{\mu\nu}=\mathrm{diag}(-,+,+,+)$, so that $k^{2}=-(k^{0})^{2}+|\vec{k}|^{2}$. Timelike momenta satisfy $k^{2}<0$, while spacelike momenta satisfy $k^{2}>0$. We work in natural units $\hbar=c=1$ and $8\pi G=M_{p}^{-2}$.
II The dual-IHO: Berry–Keating and phase space structure
The full graviton propagator of (1) around Minkowski spacetime, in de Donder gauge, can be written with the spin projectors $\left(P^{(2)}_{\mu\nu,\rho\sigma}\right)$ (spin–2) and $\left(P^{(0\text{s})}_{\mu\nu,\rho\sigma}\right)$ (spin-0) as [21]
$D_{\mu\nu,\rho\sigma}(k)=-\frac{i}{M_{p}^{2}}\left[\frac{P^{(2)}}{k^{2}}-\frac{P^{(2)}}{k^{2}-\frac{M_{p}^{2}}{\beta}}-\frac{1}{2}\frac{P^{(0\text{s})}}{k^{2}}+\frac{1}{2}\frac{P^{(0\text{s})}}{k^{2}+m^{2}}\right].$ (2)
where
$\frac{M_{p}^{2}}{\beta}=-m_{2}^{2}\,\,\left(\text{if }\beta<0\right),\qquad\frac{M_{p}^{2}}{\beta}=\mu_{2}^{2}\,\,\left(\text{if }\beta>0\right),\qquad m^{2}=\frac{M_{p}^{2}}{6\alpha}.$ (3)
where $m^{2}>0$ is mass square of the scalaron, and $m_{2}^{2}>0$ (or $\mu_{2}^{2}>0$) is the mass scale associated with the additional spin-2 mode of the theory. The theory has a massless spin-2 and a massive scalar (for $\alpha>0$) (i.e., scalaron) as an on-shell (also off-shell) propagating degree of freedom. The status of additional spin-2 depends on the coefficient of the Weyl square. If $\beta<0$, the spin-2 becomes a pathological ghost and can be an on-shell mode, whereas in the $\beta>0$ case, it becomes a dual IHO spin-2 field, which is a purely off-shell mode as we explain below. The four systems relevant to quadratic gravity are distinguished by their Hamiltonians:
$H_{\rm HO}$ $=\tfrac{\omega}{2}(\tilde{p}^{2}+\tilde{q}^{2}),\quad\text{positive definite},$ (4)
$H_{\rm ghost}$ $=-\tfrac{\omega}{2}(\tilde{p}^{2}+\tilde{q}^{2}),\quad\text{negative definite},$ (5)
$H_{\mathrm{IHO}}$ $=\tfrac{\omega}{2}(\tilde{p}^{2}-\tilde{q}^{2}),\quad\text{indefinite},$ (6)
$H_{\mathrm{dIHO}}$ $=\tfrac{\omega}{2}(-\tilde{p}^{2}+\tilde{q}^{2}),\quad\text{indefinite}.$ (7)
The IHO and dual-IHO Hamiltonians are indefinite, not negative definite. This single fact separates them categorically from the ghost. In the canonically rotated coordinates $Q=(\tilde{p}+\tilde{q})/\sqrt{2}$, $P=(\tilde{p}-\tilde{q})/\sqrt{2}$, both take the Berry–Keating (BK) form [4]:
$H_{\mathrm{IHO}}=\tfrac{\omega}{2}(QP+PQ),\qquad H_{\mathrm{dIHO}}=-\tfrac{\omega}{2}(QP+PQ).$ (8)
The BK Hamiltonian $H=xp$ is famous for its connection to the Riemann hypothesis [4]: its energy spectrum, under appropriate quantization conditions, reproduces the non-trivial zeros of the Riemann zeta function. Both the IHO and dual-IHO share hyperbolic classical trajectories $Q=Q_{0}e^{\omega t}$, $P=P_{0}e^{-\omega t}$, and the same phase-space structure up to a 90-degree canonical rotation (Fig. 1). The dual-IHO phase space is an equivalent physical system with the roles of positive and negative energy regions swapped.
Figure 1: Phase space of the IHO (left, $H_{\mathrm{IHO}}=\frac{\omega}{2}(\tilde{p}^{2}-\tilde{q}^{2})$) and the dual-IHO (right, $H_{\mathrm{dIHO}}=\frac{\omega}{2}(-\tilde{p}^{2}+\tilde{q}^{2})$). Both are related by a 90-degree canonical rotation and share hyperbolic phase-space trajectories. The separatrices divide each phase space into four regions with two distinct arrows of time, requiring a direct-sum quantization framework [16, 15, 14]. Neither system is a ghost.
The quantization of the IHO/dual-IHO poses a fundamental challenge, namely the fact that phase-space regions with opposite arrows of time (separated by the separatrices in Fig. 1) cannot be consistently handled by a single Schrödinger equation. This is precisely the open problem identified by Berry and Keating [4]. The resolution is provided by direct-sum quantum field theory (DQFT) [16, 15, 14, 13, 12, 7], which quantizes the IHO on a split Hilbert space $\mathcal{H}=\mathcal{H}_{+}\oplus\mathcal{H}_{-}$ with geometric superselection sectors corresponding to parity conjugate regions. The BK antipodal identification $(Q,P)\sim(-Q,-P)$ then acquires a geometric interpretation via the superselection structure. In this paper, we use only the consequence that the dual-IHO admits no global Fock vacuum. For quadratic gravity, with $\beta>0$, the transverse and traceless (TT) spin-2 sector after diagonalization gives the Hamiltonian [16]
$H_{\rm TT}=\frac{1}{2}\sum_{s}\int\!\frac{d^{3}k}{(2\pi)^{3}}\bigl(\dot{u}_{s}^{2}+k^{2}u_{s}^{2}\bigr)-\frac{1}{2}\sum_{s}\int\!\frac{d^{3}k}{(2\pi)^{3}}\bigl(\dot{v}_{s}^{2}+(k^{2}-\mu_{2}^{2})v_{s}^{2}\bigr),$ (9)
where $\mu_{2}^{2}\equiv M_{P}^{2}/\beta>0$. The field $u_{s}$ is the massless graviton and each mode of $v_{s}$ takes the form
$H_{s,\lambda}=-\tfrac{1}{2}p_{s,\lambda}^{2}+\tfrac{1}{2}(-\lambda+\mu_{2}^{2})q_{s,\lambda}^{2},$ (10)
where
$v_{s}(t,\vec{x})=\sum_{\lambda}q_{s,\lambda}(t)f_{\lambda}(\vec{x}),\qquad\pi_{s}(t,\vec{x})=-\dot{v}_{s}=-\sum_{\lambda}p_{s,\lambda}(t)f_{\lambda}(\vec{x}).$ (11)
with
$\nabla^{2}f_{\lambda}(\vec{x})=\lambda\,f_{\lambda}(\vec{x}),\qquad\lambda\geq 0,$ (12)
Notice that the Hamiltonian (10) is exactly a dual-IHO system. Moreover, $v_{s}$ is a dual-IHO field, which can be understood as an infinite collection of dual-IHOs, with no global Fock vacuum.
III Källén–Lehmann representation with a vanishing spectral density
The full nonperturbative Feynman propagator of any scalar field admits the Källén–Lehmann (KL) representation [11, 18]
$\widetilde{\Delta}(k^{2})=\frac{-i}{k^{2}+m^{2}-i\epsilon}+\int_{0}^{\infty}\!ds\,\frac{-i\rho(s)}{k^{2}+s-i\epsilon},$ (13)
with spectral density $\rho(s)=\sum_{n}|\langle k,n|\phi(0)|0\rangle|^{2}\delta(s-M_{n}^{2})\geq 0$. The lower limit $s=M_{n}^{2}\geq 0$ is not an independent assumption. It is a direct consequence of the Wightman spectrum condition  [22, 8]:
$\mathrm{spec}(\hat{P}^{\mu})\subset\bar{V}^{+}=\{p^{\mu}:\,p^{2}\leq 0,\;p^{0}\geq 0\},$ (14)
which requires every physical state $|n\rangle$ to satisfy $M_{n}^{2}=-p_{n}^{2}\geq 0$. The IHO and dual IHO fields with spacelike poles are excluded from the KL domain a priori by the spectrum condition. Due to this reason, the studies of QFT with tachyons in [9, 10] are not applicable in our consideration. Therefore, since the spectral parameter $s$ is restricted to non-negative values, the vanishing of the spectral density for the dual-IHO spin–2 field, $\rho(s)=0$, follows from two independent reasons.
Reason 1 (absence of a normalizable vacuum): The KL matrix element $\langle 0|\phi(0)|n\rangle$ requires a normalizable bra $\langle 0|$ with $\langle 0|0\rangle=1$. The eigenvalue equation of the dual-IHO,
$\tfrac{\omega}{2}(\tilde{q}^{2}+\partial_{\tilde{q}}^{2})\psi_{E}=E\psi_{E},$ (15)
[…]
$D_{\nu}(\tilde{q})\sim\tilde{q}^{\nu}e^{-\tilde{q}^{2}/4}+\frac{\Gamma(-\nu)}{\sqrt{2\pi}}\tilde{q}^{-\nu-1}e^{+\tilde{q}^{2}/4},\quad\tilde{q}\to-\infty$, (16)
show that the second branch grows as $e^{+\tilde{q}^{2}/4}$: not square-integrable for any real energy $E$. No eigenstate of the dual-IHO Hamiltonian belongs to $L^{2}(\mathbb{R})$; no normalizable vacuum exists. The KL matrix element is undefined, and one must set $\rho(s)=0$.
Reason 2 (spacelike pole): The field equation of the $v_{s}$ sector, $\ddot{v}_{s}-\nabla^{2}v_{s}-\mu_{2}^{2}v_{s}=0$, gives a propagator pole at
$k^{2}=\mu_{2}^{2}>0\quad\text{(spacelike in $-{+}{+}{+}$)}.$ (17)
In the KL integral (13) this would require $s=-\mu_{2}^{2}<0$, outside the spectral domain $s\geq 0$. It would correspond to a physical state with imaginary invariant mass, in direct violation of the spectrum condition (14). This is inadmissible. Both reasons independently enforce $\rho(\mu_{2}^{2})=0$. The $\delta(k^{2}-\mu_{2}^{2})$ piece is absent identically from the propagator, not suppressed by kinematics, not removed by prescription, but absent as a theorem. The propagator retains only its dispersive part,
$\Delta_{\mathrm{dIHO}}(k)=\mathrm{PV}\!\left(\frac{i}{k^{2}-\mu_{2}^{2}}\right).$ (18)
This is the unique propagator compatible with the KL representation and the Hilbert-space structure of the dual-IHO sector. It is not a fakeon projection, not a Lee–Wick contour, and not a Feynman–Wheeler prescription. Let us notice that both the Feynman propagator $F=-i/(k^{2}-\mu_{2}^{2}-i\epsilon)$ and the anti-Feynman propagator $\bar{F}=i/(k^{2}-\mu_{2}^{2}+i\epsilon)$ require $\rho\neq 0$ to be defined in the KL representation. Here, in the context of dual IHO spin-2 field, $\rho=0$ is consequence of the lack of a normalized ground state and the spacelike nature of the pole. One cannot define a vacuum for the dual IHO spin-2 field, this is completely an off-shell degree of freedom participating in the virtual interactions of massless gravitation field of the quadratic gravity. Therefore, the identity $\mathrm{PV}=\frac{1}{2}(F+\bar{F})$ (a distributional identity which is algebraically valid) is physically not possible to define for the dual-IHO spin-2 field. This is the precise reason why the objections of [2, 9] about timelike momentum supported PV propagators do not apply to the present case. We note that recent work on tachyons [10] with PV propagator ignores the subtleties with the KL representation associated with IHO and dual IHO fields. Table 1 summarizes the key distinctions.
System Hamiltonian Ground state Pole $\quad\rho$
HO pos. def. $L^{2}$ yes timelike $k^{2}<0$ $\quad\neq 0$
Ghost neg. def. $L^{2}$ yes timelike $k^{2}<0$ $\quad\neq 0$
IHO indefinite $L^{2}$ no spacelike $k^{2}>0$ $\quad=0$
dual-IHO indefinite $L^{2}$ no spacelike $k^{2}>0$ $\quad=0$
Table 1: Key properties in +++-{+}{+}{+} signature. The IHO and dual-IHO share spacelike poles and no normalizable ground state; both independently force $\rho=0$.
To be more specific, if we apply the Cutkosky cutting rule to the scattering process in quadratic gravity $\beta>0$, the optical theorem $2\,\mathrm{Im}\,\mathcal{M}(a\to a)=\sum_{f}\int d\Pi_{f}|\mathcal{M}(a\to f)|^{2}$ is trivially satisfied. Since the spin-2 IHO field propagator is (18), as its propagation is for space-like momenta as such forbidden to be an on-shell state in any kinematic scattering process, and since its spectral density also vanishes, it contributes zero imaginary part to any amplitude. For the ghost ($\beta<0$, timelike pole at $k^{2}=-m_{2}^{2}<0$), the tree-level amplitude for $2\to 2$ graviton scattering gives
$2\,\mathrm{Im}\,\mathcal{M}_{\rm ghost}=-\frac{2\pi\mathcal{N}}{M_{P}^{2}}\,\delta(k^{2}+m_{2}^{2})<0,\quad\mathcal{N}>0,$ (19)
where $\mathcal{N}=V^{\mu\nu\rho\sigma}P^{(2)}_{\mu\nu\alpha\beta}V^{\alpha\beta}_{\rho\sigma}>0$ collects all real positive kinematic factors from the vertices and spin-2 projector contracted on physical external graviton polarizations. We can clearly see the spin-2 ghost is in direct violation of the optical theorem. Since the PV propagator (18) is the only propagator the dual-IHO spin-2 field admits, every amplitude in which it participates be at the tree level or at any loop order, takes the form
$\mathcal{M}_{\rm dIHO}\propto{\rm PV}\big[\frac{1}{k^{2}-\mu_{2}^{2}}\big]$ (20)
Hence, no imaginary part and no optical theorem violation. Thus, the Cutkosky rules do not place the additional spin-2 as on-shell states. In the context of $2\to 2$ forward graviton scattering, $s=(p_{1}+p_{2})^{2}<0$ (timelike), since the dual-IHO pole sits at $k^{2}=\mu_{2}^{2}>0$ (spacelike): the on-shell condition is kinematically inaccessible. Therefore, the optical theorem is saturated only by the massless graviton and the scalaron. Thus, the dual IHO spin-2 field dictates the massless graviton scatterings where it is virtually involved i.e., t,ut-,u- channels, and at all loop levels.
IV Full Landau-equation review
Having established that the dual-IHO propagator is purely dispersive, we now verify that no physical singularities arise at loop level. A Landau-equation analysis of the relevant bubble diagrams confirms that the spacelike pole structure prevents any real kinematic pinch, and that the nonlocal UV divergences found by Anselmi [2] for timelike principal-value propagators are structurally absent in our case.
A physical Landau singularity (branch point) requires all Feynman parameters $\alpha_{i}$ to be simultaneously real and non-negative. We analyze below three distinct bubble configurations, with two involving the dual-IHO respectively shown in Fig. 2.
[…]
Figure 2: Bubble diagrams representing 1-loop self energy of a propagating massless spin-2 field (external big wavy lines) with momentum $P$. The loop wavy lines represent dual-IHO PV propagators. The solid line on the right panel denotes ordinary Feynman propagator of scalar field (the massive scalaron in our case) (a) Dual-IHO bubble: spacelike branch point $P^{2}_{*}=4\mu_{2}^{2}>0$. (b) Mixed bubble: complex Landau parameters, no physical singularity.
  1. 1.
    Dual-IHO bubble (two spacelike PV lines): Both PV lines (in Fig. 2(a)) have pole at $k^{2}=\mu_{2}^{2}>0$. The Landau on-shell conditions are $k^{2}=\mu_{2}^{2}$ and $(P-k)^{2}=\mu_{2}^{2}$. Subtracting: $P\cdot k=P^{2}/2$. Writing $k^{\mu}=aP^{\mu}+k_{\perp}^{\mu}$ with $P\cdot k_{\perp}=0$ gives $a=1/2$, then
    $k_{\perp}^{2}=\mu_{2}^{2}-\frac{P^{2}}{4}.$ (21)
    The branch point (pinch at $k_{\perp}^{\mu}=0$) is at
    $P^{2}_{*}=4\mu_{2}^{2}>0\quad\text{(which is spacelike).}$ (22)
    For physical timelike $P^{2}<0$: $k_{\perp}^{2}=\mu_{2}^{2}-P^{2}/4>\mu_{2}^{2}>0$ always. The simultaneous on-shell conditions cannot be satisfied and no absorptive part is generated.
  2. 2.
    Mixed bubble (one timelike Feynman, one spacelike PV). Let us consider here a one-loop with a Feynman propagator of the scalaron, with a pole at $k^{2}=-m^{2}<0$, and a dual-IHO spin-2 PV with a pole at $k^{2}=\mu_{2}^{2}>0$ (See Fig. 2(b)). The Landau equations of this physical process can be written as
    $\alpha_{1}(k^{2}+m^{2})=0,\quad\alpha_{2}\bigl((P-k)^{2}-\mu_{2}^{2}\bigr)=0,\quad\alpha_{1}k^{\mu}=\alpha_{2}(P-k)^{\mu}.$ (23)
    The momentum equation gives $k^{\mu}=\frac{\alpha_{2}}{\alpha_{1}+\alpha_{2}}P^{\mu}$, so
    $\left(\frac{\alpha_{1}}{\alpha_{1}+\alpha_{2}}\right)^{\!2}P^{2}=-m^{2},\quad\left(\frac{\alpha_{2}}{\alpha_{1}+\alpha_{2}}\right)^{\!2}P^{2}=\mu_{2}^{2}.$ (24)
    Dividing:
    $\frac{\alpha_{2}^{2}}{\alpha_{1}^{2}}=-\frac{\mu_{2}^{2}}{m^{2}}<0\;\Longrightarrow\;\frac{\alpha_{2}}{\alpha_{1}}=i\frac{\mu_{2}}{m}\quad(\frac{\mu_{2}}{m}\in\mathbb{R}).$ (25)
    The ratio is purely imaginary. No real non-negative solution exists for any real $P^{2}$. The threshold lies in the complex $P^{2}$ plane; therefore, no physical Landau singularity and no nonlocal divergence are present.
  3. 3.
    Feynman–Wheeler mixed bubble (Anselmi [2]): Here we consider the mixed bubble diagram (Fig. 2(b)) with the notation used in [2] with the context of ghost where both lines have timelike poles at $k^{2}=-m_{i}^{2}<0$ (one Feynman, one with the PV/FW prescription). The same Landau analysis gives
    $\left(\frac{\alpha_{i}}{\alpha_{1}+\alpha_{2}}\right)^{\!2}P^{2}=-m_{i}^{2}<0,\quad i=1,2,$ (26)
    from which
    $\frac{\alpha_{2}}{\alpha_{1}}=\frac{m_{2}}{m_{1}}>0\quad\text{(real and positive).}$ (27)
    A physical Landau singularity exists at the real timelike threshold $P^{2}=-(m_{1}\pm m_{2})^{2}<0$, accessible for physical external momenta. This is the origin of the nonlocal UV divergence found in [2],
    $\Sigma^{\prime}_{\rm div}(P)\propto\frac{m_{1}^{2}-m_{2}^{2}}{P^{2}}\,\ln\Lambda^{2},$ (28)
    which cannot be absorbed by a local counterterm and destroys renormalizability. The physical mechanism is clear. Anselmi’s nonlocal divergence requires the coexistence of $i\epsilon$ and $i\epsilon$ at timelike poles in the same loop. When one pole is spacelike, eq. (24) gives opposite signs on the right-hand sides, forcing $\alpha_{2}/\alpha_{1}$ into the complex plane. No real configuration puts both lines on-shell simultaneously. Table 2 summarizes all three cases.
Bubble Pole types $\alpha_{2}/\alpha_{1}$ Threshold
Dual-IHO [this work] both spacelike $1\in\mathbb{R}^{+}$ $P^{2}_{*}=4\mu_{2}^{2}>0$ (spacelike)
Mixed [this work] one each $i\mu_{2}/m\in i\mathbb{R}$ complex (no physical pinch)
FW [2] both timelike $m_{2}/m_{1}\in\mathbb{R}^{+}$ $P^{2}=-(m_{1}{\pm}m_{2})^{2}<0$, nonlocal div.
Table 2: Landau-equation comparison. Nonlocal divergences arise only when both poles are timelike (FW case) and $\alpha_{2}/\alpha_{1}$ is real positive, giving a physical threshold. Spacelike poles drive the ratio complex, eliminating any physical pinch.
IV.1 Locality of counterterms
The UV behavior of the dual-IHO propagator matches an ordinary Feynman propagator:
$\mathrm{PV}\!\left(\frac{1}{k^{2}-\mu_{2}^{2}}\right)=\frac{1}{k^{2}}\!\left(1+\frac{\mu_{2}^{2}}{k^{2}}+\mathcal{O}(k^{-4})\right),\quad|k^{2}|\to\infty.$ (29)
The difference between the Feynman and PV propagation is
$\frac{1}{k^{2}-\mu_{2}^{2}-i\epsilon}-\mathrm{PV}\!\left(\frac{1}{k^{2}-\mu_{2}^{2}}\right)=i\pi\delta(k^{2}-\mu_{2}^{2}),$ (30)
on the finite on-shell hypersurface $k^{2}=\mu_{2}^{2}$ at spacelike momenta. UV divergences are therefore identical in both cases. If one follows the Feynman propagator, it gives local counterterms by Stelle’s power-counting argument [21], so does our PV propagator of the dual IHO spin-2. The on-shell term $i\pi\delta(k^{2}-\mu_{2}^{2})$ in (30) is precisely the cut-producing piece that is absent because $\rho=0$.
The full spin-2 propagator reads [16]
$D^{(2)}_{\mu\nu\rho\sigma}(k)=\frac{-i}{M_{p}^{2}}\,P^{(2)}_{\mu\nu\rho\sigma}\left[\frac{1}{k^{2}-i\epsilon}-\mathrm{PV}\!\left(\frac{1}{k^{2}-\mu_{2}^{2}}\right)\right]\sim\frac{\mu_{2}^{2}}{k^{4}},\quad|k^{2}|\to\infty,$ (31)
retaining the $1/k^{4}$ UV falloff responsible for power-counting renormalizability. The dual-IHO acts as a Pauli–Villars-like subtraction: UV improvement without physical cuts.
IV.2 Comparison with existing approaches
For $\beta<0$ (ghost), the timelike pole carries a wrong-sign residue with $\rho\neq 0$: the Feynman propagator is admissible but its imaginary part violates the optical theorem. The fakeon [1, 19] removes the ghost from asymptotic states by average continuation from Euclidean space, abandoning analyticity above threshold and also leading to microcausality violation. The Lee–Wick prescription [17, 6] modifies the $k^{0}$ contour or interprets the ghost as an unstable resonance, and the PT-symmetric quantization [3] redefines the inner product. All the mentioned prescriptions start from a timelike pole and ask how to virtualize it. For $\beta>0$ (dual-IHO), the pole is spacelike and consequently $\rho=0$. There is no Feynman propagator to start from, no ghost to remove, no contour to deform. The PV propagator (18) is the only propagator the theory admits (due to the KL representation failure), and its cut vanishes, not by construction but by necessity.
V Conclusions
Quadratic gravity is the unique perturbatively renormalizable local theory of quantum gravity in four spacetime dimensions, but its extra spin–2 sector has long been regarded as the source of non-unitarity. The result of this paper is that this conclusion is not inevitable. For a positive Weyl-squared coefficient, the additional spin–2 mode is not a ghost associated with a negative definite Hamiltonian but a dual inverted harmonic oscillator (IHO) associated with an indefinite Hamiltonian (equivalent to IHO). It is worth noting that IHO is a universal physical system that appears across different areas of physics [23, 24, 7]. Starting with the dual inverted harmonic oscillator (IHO) condition ($\beta>0$), the additional spin-2 sector is characterized by a non-normalizable ground state and the presence of a spacelike pole. These two facts force the spectral density of the dual IHO spin-2 field $\rho$ to vanish entirely, which in turn reduces the propagator $\Delta$ to its PV part. Ultimately, this implies that the theory lacks the necessary branch cuts associated with physical particle production which indicate the dual IHO spin-2 cannot have any particle interpretation, no vacuum, not an asymptotic state but dictates the virtual physical processes. The main results obtained here can be summarized as:
  • For a dual IHO spin–2 field the Wightman spectrum condition requires all physical states to have real non-negative mass. A spacelike pole $k^{2}=\mu_{2}^{2}>0$ would require $s=-\mu_{2}^{2}<0$ in the KL integral, which is impossible. Independently, no normalizable vacuum exists. Both conditions, independently, force the spectral density to vanish $\rho=0$.
  • The principal value propagator is unique. Eq. (18) is the only KL consistent propagator for the dual-IHO. The identity $\mathrm{PV}=\frac{1}{2}(F+\bar{F})$ is algebraically valid but is empty of physical meaning, since neither the Feynman or anti-Feynman propagators are allowed for $\rho=0$.
  • The dual-IHO bubble has threshold $P^{2}_{*}=4\mu_{2}^{2}>0$ (spacelike), kinematically inaccessible for physical processes (Eq. (22)).
  • The mixed bubble (one timelike, one spacelike pole) gives $\alpha_{2}/\alpha_{1}=i\mu_{2}/m$ (which is purely imaginary, see Eq. (25)), therefore there is no real kinematic pinch and hence no nonlocal divergence. This is the decisive structural difference from the Feynman–Wheeler case [2] (Table 2).
  • The Feynman–PV propagator difference (30) has no UV support, locality of counterterms and the $1/k^{4}$ scaling (31) remain intact. Therefore, renormalizability preserved.
Unitarity and renormalizability coexist in quadratic gravity for $\beta>0$ without any prescription beyond the structure of the theory. The DQFT quantization framework underlying this construction, including direct-sum Hilbert spaces, geometric superselection sectors, and the resolution of the two-arrows-of-time problem of the IHO identified by Berry and Keating, is developed in Refs. [16, 15, 14, 13, 7], together with cosmological implications for inflationary gravitational waves and CMB parity asymmetry. Thus, we have a unique unitary quantum gravity in 4D that is predictive and testable with future observations.
Acknowledgements.
K.S.K. acknowledges support from the Royal Society Newton International Fellowship. Research funded by FCT grant UIDB/MAT/00212/2025 and COST action 23130. We thank David Wands for the positive encouragement and suggestion to write this short paper. We also thank Marco Piva, Krzysztof Jodlowski for the useful comments on our companion paper [16] related to the aspects of PV propagator. We also thank Luca Buoninfante for useful discussions over the years.